# Agnostic Smoothed Online Learning

### v2

@@ 5.3 Bounding the regret term for each depth
which satisfies $w\geq 2$. We combine Lemmas 12 and 13 both for the probability tolerance $\delta$, which implies that for any $w\geq 2$ and $q\in(0,1]$ satisfying Eq. 19 for $r=2$, with probability at least $1-2\delta$,
@@
Using Eqs. 20 and 21, we can now bound the regret terms from the decomposition in Eq. 14. We start with the term involving the quantities $\Delta_{k}^{(p)}$. For $p\in\{p_{0},\ldots,P\}$ we let
@@
As a result, on the event $\mathcal{E}_{\delta}$, we can combine Eqs. 22 and 23 to obtain
@@
We now combine the bounds from Eqs. 24 and 25 within the regret decomposition from Eq. 14 which shows that on $\mathcal{E}_{\delta}\cap\mathcal{F}_{\delta}$ of probability at least $1-2c\delta\ln^{2}T$,
@@ 5.4 From oblivious to adaptive benchmarks for classification
We now put the two previous estimate with Sections 5.4 and 29, for $k=\frac{1}{\sigma}\ln\frac{T}{\delta}$ and $\epsilon=1/(Tk)$. We note that the bound from Section 5.4 is vacuous if $\frac{1}{\sigma}\ln\frac{T}{\delta}=k\gtrsim T$. Hence, without loss of generality, we can suppose $k\lesssim T$. Similarly, we can suppose that $\ln\frac{T}{\delta}\lesssim\sigma T$. Altogether, this shows that with probability at least $1-2\delta$, we still have
@@ 5.5 Proof of Lemma 13
In $(i)$ we used Eq. 36 and in $(ii)$ we used Eqs. 35 and 32. Now note that conditionally on $a\leq A$ and $\mathcal{H}_{t(a)-1}$, we have
@@ Appendix B Learning with expert advice guarantee for A-Exp
On period $(T_{k-1},T_{k}]$, A-Exp exactly implements the exponentially weighted forecaster with parameter $\eta_{k}=\sqrt{2\ln K/(\Delta_{max,k}+1)}$. Hence, we can use Eq. 1 to bound the regret accumulated on this period which gives for all $T\in(T_{k-1},T_{k}]$,
@@
$PReg(T)\leq\sum_{k^{\prime}\leq k}2^{k}\sqrt{\ln K}\leq 2^{k+1}\sqrt{\ln K}\overset{(ii)}{\leq}8\sqrt{\Delta_{T}\ln K}.$
@@
If $k=1$, we have directly $PReg(T)\leq 2\sqrt{\ln K}$. This ends the proof for the bound on the pseudo-regret.
To obtain high-probability bounds on the regret $Reg(T)$, we could simply use Azuma-Hoeffding’s inequality. However, this would add an additional term $\sqrt{T\ln\frac{1}{\delta}}$ that is prohibitive for our bounds: potentially we have $\Delta_{T}\ll T$. Instead, we use Freedman’s inequality which gives a more precise control on tail probabilities for martingales. Lemma 19 applied with $Z_{t}=r_{t,\hat{i}_{t}}^{2}-\mathbb{E}_{\hat{i}_{t}}[r_{t,\hat{i}_{t}}^{2}\mid\mathcal{H}_{t}]$ for $t\in[T]$ and $\eta=1/2$ implies that with probability at least $1-\delta$,
$Reg(T)\overset{(i)}{\leq}PReg(T)+\frac{1}{2}\sum_{t=1}^{T}\mathbb{E}[r_{t,\hat{i}_{t}}^{4}\mid\mathcal{H}_{t}]+2\ln\frac{1}{\delta}\overset{(ii)}{\leq}\frac{3}{2}PReg(T)+2\ln\frac{1}{\delta},$
@@ Appendix D Proofs from Section 4
In this section, we prove the results related to the simplified algorithm Cover. These are essentially simplified proofs of their counterparts for the main proofs from Section 5, hence we will only highlight the main differences.
@@
We are now ready to prove the main regret bound for Cover.
@@
Appendix E Concentration inequalities and technical lemmas

### v3

@@ 5.3 Bounding the regret term for each depth
which satisfies $w\geq 2$. We combine Lemmas 10 and 11 both for the probability tolerance $\delta$, which implies that for any $w\geq 2$ and $q\in(0,1]$ satisfying Eq. 20 for $r=2$, with probability at least $1-2\delta$,
@@
Using Eqs. 21 and 22, we can now bound the regret terms from the decomposition in Eq. 15. We start with the term involving the quantities $\Delta_{k}^{(p)}$. For $p\in\{p_{0},\ldots,P\}$ we let
@@
As a result, on the event $\mathcal{E}_{\delta}$, we can combine Eqs. 23 and 24 to obtain
@@
We now combine the bounds from Eqs. 25 and 26 within the regret decomposition from Eq. 15 which shows that on $\mathcal{E}_{\delta}\cap\mathcal{F}_{\delta}$ of probability at least $1-2c\delta\ln^{2}T$,
@@ 5.4 From oblivious to adaptive benchmarks for classification
We now put the two previous estimate with Sections 5.4 and 30, for $k=\frac{1}{\sigma}\ln\frac{T}{\delta}$ and $\epsilon=1/(Tk)$. We note that the bound from Section 5.4 is vacuous if $\frac{1}{\sigma}\ln\frac{T}{\delta}=k\gtrsim T$. Hence, without loss of generality, we can suppose $k\lesssim T$. Similarly, we can suppose that $\ln\frac{T}{\delta}\lesssim\sigma T$. Altogether, this shows that with probability at least $1-2\delta$, we still have
@@ 5.5 Proof of Lemma 11
In $(i)$ we used Eq. 37 and in $(ii)$ we used Eqs. 36 and 33. Now note that conditionally on $a\leq A$ and $\mathcal{H}_{t(a)-1}$, we have
@@ Appendix B Learning with expert advice guarantee for A-Exp
On period $(T_{k-1},T_{k}]$, A-Exp exactly implements the exponential weights forecaster with parameter $\eta_{k}=\sqrt{2\ln K/(\Delta_{max,k}+1)}$. Hence, we can use Eq. 2 to bound the regret accumulated on this period which gives for all $T\in(T_{k-1},T_{k}]$,
@@
$\textnormal{PReg}(T)\leq\sum_{k^{\prime}\leq k}2^{k}\sqrt{\ln K}\leq 2^{k+1}\sqrt{\ln K}\overset{(ii)}{\leq}8\sqrt{\Delta_{T}\ln K}.$
@@
If $k=1$, we have directly $\textnormal{PReg}(T)\leq 2\sqrt{\ln K}$. This ends the proof for the bound on the pseudo-regret.
To obtain high-probability bounds on the regret $\textnormal{Reg}(T)$, we could simply use Azuma-Hoeffding’s inequality. However, this would add an additional term $\sqrt{T\ln\frac{1}{\delta}}$ that is prohibitive for our bounds: potentially we have $\Delta_{T}\ll T$. Instead, we use Freedman’s inequality which gives a more precise control on tail probabilities for martingales. Lemma 20 applied with $Z_{t}=r_{t,\hat{i}_{t}}^{2}-\mathbb{E}_{\hat{i}_{t}}[r_{t,\hat{i}_{t}}^{2}\mid\mathcal{H}_{t}]$ for $t\in[T]$ and $\eta=1/2$ implies that with probability at least $1-\delta$,
$\textnormal{Reg}(T)\overset{(i)}{\leq}\textnormal{PReg}(T)+\frac{1}{2}\sum_{t=1}^{T}\mathbb{E}[r_{t,\hat{i}_{t}}^{4}\mid\mathcal{H}_{t}]+2\ln\frac{1}{\delta}\overset{(ii)}{\leq}\frac{3}{2}\textnormal{PReg}(T)+2\ln\frac{1}{\delta},$
@@ Appendix D Proofs from Section 4
In this section, we prove the results related to the simplified algorithm Cover. We start by giving a detailed proof sketch for Proposition 7.
Proof sketch of Proposition 7.
Fix a fixed threshold $q$. For convenience, we denote by $E_{k}=(T_{k-1},T_{k}]$ the $k$-th epoch for $k\in[K]$. We also fix a value $w\asymp d\ln\frac{T}{\sigma}$ which will play the role of $w(T,\delta)$. Given the desired result, we will only focus on times $t\in E_{k}$ for $k\in[K]$, for which $\gamma_{T_{k-1}}(t)\geq q$. For simplicity, we now assume that we have discarded all other times, so that for all $t\in E_{k}$ one has $\gamma_{T_{k-1}}(t)\geq q$. Further, note that the quantity of interest for Proposition 7 only involves epochs $k\in[K]$ for which $\sum_{t\in E_{k}}\gamma_{T_{k-1}}(t)\geq w$. Therefore, for each epoch $E_{k}$ such that $\sum_{t\in E_{k}}\gamma_{T_{k-1}}(t)\geq w$, we also discard all times $t\in E_{k}$ with $t>t_{max,k}$, where $t_{max,k}\in E_{k}$ is the first time for which $\sum_{t\in E_{k},t\leq t_{max,k}}\gamma_{T_{k-1}}(t)\geq w$. Up to renaming the remaining times in the sequence, we now consider that the epochs containing remaining times correspond to $E_{k}=(t_{k-1},t_{k}]$ for $k\in[K]$ where $t_{0}\leq t_{1}\leq\cdots\leq t_{K}$. The remaining sequence therefore has horizon $t_{K}\leq T$ (within the formal proof this is denoted by $A$). We can in fact check that these times can be deleted online, and as a result, the corresponding subsequence of $(x_{t})_{t\in[T]}$ is still $\sigma$-smooth.
Clearly, we have
|$\left|\left\{k\in[K]:\sum_{t\in E_{k}}\gamma_{t}\geq w\right\}\right|\leq\frac{1}{w}\sum_{t=1}^{t_{K}}\gamma_{t},$ (46)
where we used the shorthand $\gamma_{t}$ to denote $\gamma_{T_{k-1}}(t)$ where $t\in E_{k}$. Therefore, in the rest of the proof, we aim to bound the right-hand side $\sum_{t=1}^{t_{K}}\gamma_{t}$.
To each time $t\in[t_{K}]$ we associate a function $g_{t}$ as follows. By construction, we recall that $\gamma_{t}\geq q$. Hence, by definition of $\gamma_{t}$, we can fix $f_{t},h_{t}\in\mathcal{F}$ such that
$\mathbb{P}_{x\sim\mu_{t}}(f_{t}(x)\neq h_{t}(x))\geq\frac{\gamma_{t}}{2},$
and such that $f_{t}(x_{s})=h_{t}(x_{s})$ for all times $s\in[t_{k-1}]$ from previous epochs. We then pose $g_{t}:x\mapsto\mathbbm{1}[f_{t}(x)\neq h_{t}(x)]$. In particular,
$\sum_{t=1}^{t_{K}}\mathbb{E}[g_{t}(x_{t})\mid\mathcal{H}_{t}]\geq\frac{1}{2}\sum_{t=1}^{t_{K}}\gamma_{t}.$
We next upper bound on the right-hand side. To do so, we use the decoupling result Lemma 21 (akin to [15, Lemma 3]), which gives
$\frac{1}{2}\sum_{t=1}^{t_{K}}\gamma_{t}\leq\sum_{t=1}^{t_{K}}\mathbb{E}[g_{t}(x_{t})\mid\mathcal{H}_{t},g_{t}]\lesssim\sqrt{\frac{t_{K}\ln t_{K}}{\sigma}\left(\ln t_{K}+\sum_{t=1}^{t_{K}}\frac{1}{t}\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t},g_{t}]\right)},$ (47)
where $(x_{t}^{\prime})_{t\leq t_{K}}$ is a tangent sequence to $(x_{t})_{t\leq t_{K}}$. We next fix $t\in E_{k}$ and bound
$\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t},g_{t}]\leq\sum_{s=1}^{t_{k-1}}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t},g_{t}]+\sum_{s=t_{k-1}+1}^{t-1}\gamma_{t}\leq\mathbb{E}_{x_{1}^{\prime},\ldots,x_{t_{K}}^{\prime}}\left[\sum_{s=1}^{t_{k-1}}g_{t}(x_{s}^{\prime})-2g_{t}(x_{s})\right]+w.$
In the last inequality, we used the fact that by construction, $t_{k}$ is the first index for which $\sum_{t\in E_{k},t\leq T_{k}}\gamma_{t}\geq w$, as well as the fact that for all $s\leq t_{k-1}$, $g_{t}(x_{s})=f_{t}(x_{s})-h_{t}(x_{s})=0$. An analog term to the right-hand side expectation was bounded in [15, Theorem 2] in terms of the Wills functional of the function class $\mathcal{F}$. For our purposes, we need to use a high-probability variant Lemma 23. After computations, we obtain with high probability,
$\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t},g_{t}]\lesssim\ln\mathbb{E}_{\mu}\left[W_{4T\ln T/\sigma}\left(\mathcal{F}\right)\right]+\ln T+w\lesssim d\ln\frac{T}{\sigma}+w\lesssim w.$
We plug this into Eq. 47 and using the fact that $\ln t_{K}\leq\ln T$ and $t_{K}\leq\frac{1}{q}\sum_{t=1}^{t_{K}}\gamma_{t}$ by construction, which gives
$\sum_{t=1}^{t_{K}}\gamma_{t}\lesssim\frac{\ln^{2}T}{q\sigma}\cdot w.$
Together with Eq. 46 this proves the desired bound.
We now formally prove Proposition 7. As a remark, how the epochs are constructed is very flexible: we considered in Proposition 7 the fixed schedule of epochs for Cover ($(T_{k})_{k\in[K]}$) but randomized epochs are also possible, which may be useful for improved regret bounds in the regression case. The proof is essentially a simplification from its counterpart Lemma 11, hence we will only highlight the main differences. We start by proving Proposition 7. In this result, how the epochs are constructed is very flexible: Proposition 7 considered fixed schedule for Cover but randomized epochs are also possible, which may be useful for improved regret bounds in the regression case. We prove the corresponding generalization below.
Proposition 17.
@@
We are now ready to prove the main regret bound for Cover, which is also essentially the same as for its counterpart within Section 5.
@@
Appendix E Technical lemmas
E.1 Covering bounds
We first state the classical Sauer-Shelah’s lemma [45, 46] which bounds the size of the projection of a function class with finite VC dimension onto a set $\{x_{1},\ldots,x_{n}\}\subset\mathcal{X}$.
Lemma 18 (Sauer-Shelah’s lemma).
Generalizing Sauer-Shelah’s lemma, it is known that the fat-shattering dimension can be used to bound the size of empirical covers regression function classes.
Theorem 19 (Theorem 4.4 from [44]).
E.2 Concentration inequalities